the area of a rectangular garden is 81 square meter. the width is 6 meters longers than one-third of the length. find the length and the width of the garden
Answer by unlockmath(1688)

Hello,
Let x represent the length.
x(1/3x+6)=81
Expand out and multiply by 3:
x^2+18x=243
Subtract 243:
x^2+18x-243=0
Factor:
(x-9)(x+27)
Solve for x:
x=9
x=-27
9 is the only sensible answer so:
length is 9 and width is 9 meters.
It's a square.
Make sense?
RJ
www.math-unlock.com